OpenStudy (anonymous):

solve for x 3x+sqrt(x-1)=x+2

OpenStudy (anonymous):

I tried it, but got something wrong, and I need someone to check it

OpenStudy (zzr0ck3r):

subtract 3x from sides, then square both sides and note \((\sqrt{x})^2=x\)

OpenStudy (anonymous):

9x^2+x-1=x^2+4 x^2+x=5 x^2+x-5=0

OpenStudy (anonymous):

so like x-1=x^2+4+9x^2

OpenStudy (anonymous):

then to make it a quadratic don't I go 0=10x^2-x+5?

OpenStudy (zzr0ck3r):

\(3x+\sqrt{x-1}=x+2\iff \sqrt{x-1}=-2x+2\iff (x-1)=(-2x+2)^2\)

OpenStudy (anonymous):

<maybe that's where I messed up> that would simplify to x-1=4x+4

OpenStudy (jhannybean):

If you subtract 3x from both sides of your equation, you can then square both sides to get rid of the square root on the left side. \[3x+\sqrt{x-1}=x+2\]\[\sqrt{x-1} = x+2 - 3x\]\[(\sqrt{x-1})^2 = (-2x +2)^2\]

OpenStudy (anonymous):

sorry *x^2 x-1=4x^2+4

OpenStudy (anonymous):

0=4x^2-x+5 but since we can't factor it, ill do the -b+/-sqrt(b^2-4ac)/2a

OpenStudy (jhannybean):

Why are you even bothering to do that?

OpenStudy (zzr0ck3r):

\(3x+\sqrt{x-1}=x+2\iff \sqrt{x-1}=-2x+2\iff (x-1)=(-2x+2)^2\) \(x-1=4x^2-8x+4 \iff 4x^2-9x+5=0\)

OpenStudy (zzr0ck3r):

this factors

OpenStudy (anonymous):

I don't understand how you did that

OpenStudy (jhannybean):

He moved everything over to the right.

OpenStudy (anonymous):

how did (-2x+2)^2 become 4x^2-8x+4?

OpenStudy (anonymous):

won't that be (-2)^2*(x)^2+(2)^2 4x^2+4

OpenStudy (jhannybean):

\[(a +b)^2 = a^2 +2ab+b^2\]

OpenStudy (anonymous):

oh, I didn't see that sorry now that makes sense then 4x^2-9x+5=0 is the quadratic answer, then factoring it (x-1)(4x-5) 1, 5/4 is the answer, just need to check it

OpenStudy (jhannybean):

mmhmm.

OpenStudy (anonymous):

thank you!

OpenStudy (jhannybean):

\[4x^2-9x+5=0\]\[x^2 -9x +20=0\]\[(x-5)(x-4)=0\]\[\left(x-\frac{5}{4}\right)\left(x-\frac{4}{4}\right)=0\]\[(4x-5)(x-1)=0\]\(\checkmark\) good.